If Steve's age is decreased by 12 years, and that difference is multiplied by 3, the result is 24 years. what is steve's age?

Answer by ReadingBoosters(3246)

work backwards
24 / 3 = 8 + 12 = 20 years old
.........
Proof:
20 - 12 = 8 * 3 = 24